OpenStudy (anonymous):

9 b + 42 = -12 Please help me solve?

OpenStudy (anonymous):

9b = (-12 - 42) 9b = (-54) b = (-54/9) b = -6

OpenStudy (anonymous):

Thanks, but now I'm confused... What type of problem is this? Multi-step Equation? b = 3.3 was my original answer because I subtracted -42 from both sides which made it "9b = 30" so then I divided 9 onto both sides and got b=3.3 But you added the 42 and -12 together then divided by 9.. Is this a different type of problem I'm dealing with? thanks

OpenStudy (phi):

because I subtracted -42 from both sides that is the correct step but you have to be careful about the signs: 9 b + 42 -42 = -12-42 the 42-42 on the left=0 but -12 - 42 is not 30!

OpenStudy (phi):

-12 - 42 means you start at -12 and go -42 more in the negative direction

OpenStudy (anonymous):

Ohh so -54 divided by 9 which equals -6 So the guy above us solved it a different way Thanks man :')

OpenStudy (anonymous):

no wait , he actually solved it the same way haha

OpenStudy (phi):

Yes, just threw in some parentheses, which do not really change things.